OpenStudy (anonymous):

Determine the largest acute angle of a right triangle which has legs of length 2 and 3

OpenStudy (anonymous):

 sin, cos, tan, 180, you have so many options. Lets solve for angle x. Opposite to x is 3, and Adjacent is 2. therefore, we should use TAN \[\tan \theta= opposite / adjacent\]\[\tan \theta= 3/2\]\[\theta=\tan^{-1} (3/2)\]\[\theta=56.3\] so now we have one angle. its easier now to find y, cause we know all angles in a triangle add up to 180. 180=90+56.3+y 180-90-56.3=y y=33.7 Our two angles are 56.3 and 33.7. Which ones larger?